In a triangle, the measure of the first angle is twice of the second angle. The measure of the third angle is 76 degrees mor than the second angle. Use the fact that the sum of the measures of the three angles of the triangle is 180 degrees to find the measure of each angle.

sm angle=
md angle=
lg angle=

Answer by stanbon(75887)

In a triangle, the measure of the first angle is twice of the second angle. The measure of the third angle is 76 degrees mor than the second angle. Use the fact that the sum of the measures of the three angles of the triangle is 180 degrees to find the measure of each angle.
-------------------
f = 2s
t = s+76
f + s + t = 180
-----
Substitute for f and t and solve for "s":
s+76 + s + 2s = 180
4s = 104
s = 26 degrees
f = 2s = 52 degrees
t = s+76 = 102
---------------------
sm angle= 26
md angle= 52
lg angle= 102
=====================
cheers,
Stan H.